OpenStudy (anonymous):

Which of the following is a solution for the system y = -3x + 5 and y = 2x + 5?

OpenStudy (anonymous):

Answers (1, 7) (5, 0) (2, -1) (0, 5)

OpenStudy (akashdeepdeb):

Okay so here we have y = -3x + 5 y = 2x + 5 Now if y equals both of them then both equal themselves! :D So -3x + 5 = 2x + 5 Now find the value of x first! :D And then put the value of x into anyone of the equations to get the value of y! :D Understood? :)

OpenStudy (anonymous):

i don't really understand, I've read your explanation about 5 times lol...

OpenStudy (akashdeepdeb):

Hahaha....No Probs! :D Okay look we are given y = -3x + 5 y = 2x + 5 So that means if a=b a=c Then b=c ['a' forms a link] So if y = -3x + 5 y = 2x + 5 Then -3x + 5 = 2x + 5 5x = 5-5 x = 0/5 = 0 Substitute this x's value in the [Let's say] The first equation.. y = -3x + 5 = -3(0) + 5 = 0+5 = 5 So x=0,y=5 Their intersection point would be (x,y) = (0,5) Getting anywhere? :D

OpenStudy (anonymous):

O.o iiiiii see now, thanks a lot!

OpenStudy (akashdeepdeb):

:)